Solve the PDE p+q=1
1
Expert's answer
2021-08-01T10:20:51-0400

Given Partial Differential equation is -


"=p+q=1"


Let "p=\dfrac{\partial u}{\partial t}" ", q=\dfrac{\partial u}{\partial x}"


"=" "\dfrac{\partial u}{\partial t}+\dfrac{\partial u}{\partial x}=1"


This gives the system of ODE's as -


"=\dfrac{dx}{1}=\dfrac{dt}{1}=\dfrac{du}{1}"


which has corresponding equation equal to some constants from there integrals-


"=x-t=c_1"


"=u-x=c_2"


Which gives ,


"u(x,t)=x+{\phi}(x-t)" , which is the required solution, where "\phi" is an arbitrary function.